# KeyMPs: One-Shot Vision-Language Guided Motion Generation by Sequencing DMPs for Occlusion-Rich Tasks

### v1

@@ Abstract
Dynamic Movement Primitives (DMPs) provide a flexible framework wherein smooth robotic motions are encoded into modular parameters. However, they face challenges in integrating multimodal inputs commonly used in robotics like vision and language into their framework. To fully maximize DMPs’ potential, enabling them to handle multimodal inputs is essential. In addition, we also aim to extend DMPs’ capability to handle object-focused tasks requiring one-shot complex motion generation, as observation occlusion could easily happen mid-execution in such tasks (e.g., knife occlusion in cake icing, hand occlusion in dough kneading, etc.). A promising approach is to leverage Vision-Language Models (VLMs), which process multimodal data and can grasp high-level concepts. However, they typically lack enough knowledge and capabilities to directly infer low-level motion details and instead only serve as a bridge between high-level instructions and low-level control. To address this limitation, we propose Keyword Labeled Primitive Selection and Keypoint Pairs Generation Guided Movement Primitives (KeyMPs), a framework that combines VLMs with sequencing of DMPs. KeyMPs use VLMs’ high-level reasoning capability to select a reference primitive through keyword labeled primitive selection and VLMs’ spatial awareness to generate spatial scaling parameters used for sequencing DMPs by generalizing the overall motion through keypoint pairs generation, which together enable one-shot vision-language guided motion generation that aligns with the intent expressed in the multimodal input. We validate our approach through an occlusion-rich manipulation task, specifically object cutting experiments in both simulated and real-world environments, demonstrating superior performance over other DMP-based methods that integrate VLMs support.
@@ I INTRODUCTION
Dynamic Movement Primitives (DMPs) are a powerful framework for robotic motion generation that encodes motions compactly with stability and robustness due to their foundation in dynamical systems [1, 2, 3]. DMPs enable efficient learning and reproduction of motor behaviors and offer scalability in spatial and temporal domains, which are crucial for performing diverse tasks in different environments [4, 5, 6]. The modularity and parameterization of DMPs generate motion in a flexible manner wherein motions are encoded into easily integrated parameters that enable them to adapt to different environmental conditions [7, 8]. However, DMPs have an inability to handle effectively multimodal inputs like vision and language data that are commonly used in robotics.
@@
In order to achieve this goal, two main objectives must be met, i.e., associating DMPs with language and vision inputs[10, 11, 12] and extending DMPs to tasks requiring one-shot vision-language guided complex motion generation, as current DMP-based motion generation methods [13, 14] focus on relatively simple motions. Achieving these objectives involves overcoming specific challenges, many of which stem from the limitations of deep-learning methods. In particular, associating DMP parameters with high-dimensional data requires extensive datasets for generalization, which thus limits their real-world use [13, 14] Additionally, interpreting linguistic inputs [15, 9] and generating long, complex DMP motions in one-shot [14] expand the required feature space. This issue intensifies the need for even larger datasets beyond what is already required for high-dimensional input. Addressing these challenges is crucial for executing intricate tasks based on high-level instructions and visual cues.
@@
Building on these features, we propose Keyword Labeled Primitive Selection and Keypoint Pairs Generation Guided Movement Primitives (KeyMPs), a framework that integrates VLMs into DMP-based motion generation. The framework has two main components that leverage VLMs’ strengths. The first, called Keyword Labeled Primitive Selection, uses VLMs’ natural language processing to select task-related intuitive labels mapped to DMP parameters (referred to as learned parameters) obtained via imitation learning [18, 19], thereby bridging natural language with DMPs. The second, called Keypoint Pairs Generation, exploits VLMs’ spatial awareness to generate multiple keypoint pairs. Each generated keypoint pair serves as distinct spatial scaling parameters which will be used to facilitate sequencing DMPs. We complement VLMs’ 3D capabilities [20] by integrating real-world depth cues (e.g., an object’s height, obstacle’s height), enabling a simplified 2D keypoint representation for spatial parameters that preserve key 3D details.
@@
    Demonstrating KeyMPs’ effectiveness in simulated environments through a comprehensive analysis of an occlusion-rich task, executable 3D motions that align with the intent expressed in the multimodal input, and showcasing its superior task generalization compared to deep learning-based and ablation methods.
@@ II-A Deep Learning-based DMPs Frameworks with Vision Input
Among these deep-learning-based DMP frameworks, only CIMEDNet [14]a, which generates motion by using a set of DMP parameters, and our prior work on DSDNet [21], which sequences DMPs by generating multiple sets of DMP parameters, are capable of planning complete task-oriented motion sequences from an image input. Despite this capability, the performance of both models remains severely limited by a lack of training data, which hinders their adaptability to new situations and tasks. To address these limitations, we used cross-domain knowledge through VLMs in this research to provide an understanding of visual input and essential insights to develop intricate robotic motions without requiring extra fine-tuning and that adapt to new situations.
@@ II-B LLM and VLM Integrated DMP Frameworks
Recent research has explored combining LLMs with DMPs in order to enhance robotic motion generation [19, 18]. In particular, it has demonstrated the feasibility of translating high-level language instructions into low-level motion primitives for more complex task planning. However, as tasks become increasingly detailed or require contextual information from visual inputs, relying solely on language-driven approaches can make it challenging to generate precise positional data [30] or manage different groups of DMP parameters [19, 18].
@@ III-B Sequencing Multiple DMPs in Time
For more intricate behaviors, multiple DMPs can be sequenced in time [37, 38, 39, 40], with each handling a segment of the motion. Here, let $K$ be the total number of segments, and define time boundaries $t_{0}\;<\;t_{1}\;<\;\cdots\;<\;t_{K},$ where $t_{0}$ is the start time and $t_{K}$ is the end time of the entire motion. The overall trajectory $\mathbf{Y}(t)$ is defined as a piecewise function:
@@ IV-C1 Keyword Labeled Primitive Selection
This approach requires collecting a small dataset of basis function weights for the specific primitives needed for the task. For example, in the object-cutting task shown in Fig. 2, primitives for different cutting styles such as straight, sawing, mincing, etc. must be collected individually. It should be noted that DMP imitation is performed on motions normalized to a starting position of 0 and a goal position of 1 for all coordinates to ensure proper scaling of the DMPs.
@@ IV-C2 Keypoint Pairs Generation
Algorithm 1 KeyMPs Motion Generation
1:  Parameters:
2:   $img_{env}$ Environment Visual Input
3:   $l$ Natural Language Input
4:   $h$ Object’s Height
5:  Initialize:
6:   $D$ Reference Primitive Dictionary
8:  $keyword\leftarrow\mathrm{VLM_{keyword}}(img_{obj},l)$
9:  $keypoint\ pairs\leftarrow\mathrm{VLM_{keypoints}}(img_{obj},l,pos_{global},h)$
10:  $motion\ sequence\leftarrow\mathrm{DMPMotionGen}(D,keyword,$ $keypoint\ pairs)$
11:  return  $motion\ sequence$
@@ V SIMULATION EXPERIMENT
To evaluate the effectiveness of our KeyMPs framework and address key research questions, we selected an occlusion-rich task, cutting objects with a knife, that requires the generation of complex motion in one shot, as the movement of the knife attached to the robot would occlude visual observations mid-execution. For this purpose, we created an environment for the object-cutting task in Isaac Gym [41], as shown in Fig. 5. We conducted two experiments designed to test various aspects of our framework.
Figure 5: Object-cutting environment in Isaac Gym simulation.
@@ V-A Research Questions
For the first experiment, we tested comparison methods with a more direct generation of DMP-based motion through VLMs to answer (P1). For the second experiment, we checked the contribution of each VLM-based component separately. Finally, for the third experiment, we tested the generalization performance of our method by comparing it with existing deep-learning-based DMP frameworks.
@@ V-B Experimental Setting
We performed three experiments focusing on the object-cutting task to answer our research questions. Below, we outline the task details that serve as the foundation for our investigation.
V-B1 Task Description
@@
Trained Task 1 Cabbage A single horizontal slice in the middle
@@ 9 Eggplant The object is 30 cm long, cut it vertically into 5 cm slices
Unseen Task 10 Cabbage Slice the object into 3 parts horizontally
@@ 15 Baguette The object is 45 cm long, cut it vertically into 5 cm slices
V-B2 Evaluation Method
For each task scenario in Table I, a singular, well-defined cutting line was established algorithmically based on the input prompt. For example, in Case 2 (banana bread with the prompt “I want to eat 1 slice for each day of this week, cut it vertically”), an algorithm generated 6 evenly spaced horizontal lines to simulate cutting the object into 7 slices. Similarly, in Case 4 (round pizza with the prompt “Cut it into 8 equal slices”), radial lines—comprising one horizontal, one vertical, and two diagonal lines—were created to divide the object into 8 pie slices.
These algorithmically generated ground-truth lines, together with the object’s height information, were then used to scale the cutting primitive paired to the object to finally produce the 3D ground-truth coordinates. In addition, for Experiment 3, we created training data of DMP parameters by imitating these ground-truth coordinates, thereby ensuring that the dataset reflected both the intended cutting strategies and the characteristics of the physical motion. To further account for real-world variability, these ground-truth motions were also applied to objects with slightly randomized dimensions, which ensured that the evaluation would capture the system’s ability to adapt to subtle variations in object size.
Figure 6: Evaluation visualization of the generated motion (light blue) to the ground-truth motion (red) by matching each unpaired ground-truth point (crimson) to the closest unpaired point in the generated motion (blue).
For a quantitative evaluation (as shown in Fig. 6), the ground-truth motion was first downsampled to a fixed number of points. Each unpaired point in this downsampled trajectory was then matched to the nearest unpaired point in the generated motion. This process enabled a consistent and accurate comparison between the generated and intended motions.
V-B3 Primitive Dictionary Preparation
We prepared two cutting primitives by having DMPs imitate predefined keypoints:
@@ V-D1 Objective
In this experiment, we evaluated the extent to which the structured component decomposition approach implemented in KeyMPs produces executable motions that align with the intent expressed in the multimodal input. Specifically, we compared KeyMPs against an unstructured, end-to-end VLM-driven approach that directly generates motion primitives. This comparison tested whether breaking the motion generation process into subtasks improves the capture of 3D geometric details and DMP parameters, resulting in motions that more faithfully reflect the intended outcome. Each task in Table I was executed ten times and was quantitatively evaluated using the criteria in §V-B2.
@@ V-D3 Results
Figure 14: Task-feature-extractor initialization prompt.
@@ V-E1 Objective
To address (P2), we evaluated the individual impacts of the two VLM-based components in our framework, i.e. keypoint pairs generation and keyword labeled primitive selection. This ablation study isolated each component’s effectiveness in producing executable motions aligned with the intended outcome. Each task in Table I was executed ten times and assessed using the criteria in §V-B2.
@@ V-F1 Objective
To address (P3), we determined whether KeyMPs achieves better generalization to unseen cutting tasks than deep-learning methods when training data is limited. The deep-learning baselines were trained on a restricted dataset of 90 demonstrations (10 per task for 9 tasks in Table I), simulating scenarios where collecting large-scale robotic data is impractical. In contrast, KeyMPs doesn’t require extensive task-specific training data, relying instead on structured component decomposition and vision-language priors. Each task in Table I was executed ten times, and the quantitative evaluation (§V-B2) focused on generalization performance, specifically success rates on novel objects and cutting patterns.

### v2

@@ Abstract
Dynamic Movement Primitives (DMPs) provide a flexible framework wherein smooth robotic motions are encoded into modular parameters. However, they face challenges in integrating multimodal inputs commonly used in robotics like vision and language into their framework. To fully maximize DMPs’ potential, enabling them to handle multimodal inputs is essential. In addition, we also aim to extend DMPs’ capability to handle object-focused tasks requiring one-shot complex motion generation, as observation occlusion could easily happen mid-execution in such tasks (e.g., knife occlusion in cake icing, hand occlusion in dough kneading, etc.). A promising approach is to leverage Vision-Language Models (VLMs), which process multimodal data and can grasp high-level concepts. However, they typically lack enough knowledge and capabilities to directly infer low-level motion details and instead only serve as a bridge between high-level instructions and low-level control. To address this limitation, we propose Keyword Labeled Primitive Selection and Keypoint Pairs Generation Guided Movement Primitives (KeyMPs), a framework that combines VLMs with sequencing of DMPs. KeyMPs use VLMs’ high-level reasoning capability to select a reference primitive through keyword labeled primitive selection and VLMs’ spatial awareness to generate spatial scaling parameters used for sequencing DMPs by generalizing the overall motion through keypoint pairs generation, which together enable one-shot vision-language guided motion generation that aligns with the intent expressed in the multimodal input. We validate our approach through experiments on two occlusion-rich tasks: object cutting, conducted in both simulated and real-world environments, and cake icing, performed in simulation. These evaluations demonstrate superior performance over other DMP-based methods that integrate VLM support.
@@ I INTRODUCTION
Dynamic Movement Primitives (DMPs) are a powerful framework for robotic motion generation. They encode motions compactly with stability and robustness due to their foundation in dynamical systems [1, 2, 3]. DMPs enable efficient learning and reproduction of motor behaviors and offer scalability in spatial and temporal domains, which are crucial for performing diverse tasks in different environments [4, 5, 6]. The modularity and parameterization of DMPs generate motion in a flexible manner wherein motions are encoded into easily integrated parameters that enable them to adapt to different environmental conditions [7, 8]. However, DMPs have an inability to handle effectively multimodal inputs like vision and language data that are commonly used in robotics.
@@
In order to achieve this goal, two main objectives must be met, i.e., associating DMPs with language and vision inputs[10, 11, 12] and extending DMPs to tasks requiring one-shot vision-language guided complex motion generation, as current DMP-based motion generation methods [13, 14] focus on relatively simple motions. Achieving these objectives requires overcoming challenges, particularly those from deep-learning methods. In particular, associating DMP parameters with high-dimensional data requires extensive datasets for generalization, which thus limits their real-world use [13, 14] Additionally, interpreting linguistic inputs [15, 9] and generating long, complex DMP motions in one-shot [14] expand the required feature space. This issue intensifies the need for even larger datasets beyond what is already required for high-dimensional input. Addressing these challenges is crucial for executing intricate tasks based on high-level instructions and visual cues.
@@
Building on these features, we propose Keyword Labeled Primitive Selection and Keypoint Pairs Generation Guided Movement Primitives (KeyMPs), a framework that integrates VLMs into DMP-based motion generation. The framework leverages VLMs’ strengths through two main components. The first, called Keyword Labeled Primitive Selection, uses VLMs’ natural language processing to select task-related intuitive labels mapped to DMP parameters (referred to as learned parameters) obtained via imitation learning [18, 19], thereby bridging natural language with DMPs. The second, called Keypoint Pairs Generation, exploits VLMs’ spatial awareness to generate multiple keypoint pairs. Each generated keypoint pair serves as distinct spatial scaling parameters which will be used to facilitate sequencing DMPs. We enhance VLMs’ 3D capabilities [20] by integrating real-world depth cues, such as an object’s or obstacle’s height. This allows a simplified 2D keypoint representation for spatial parameters while preserving key 3D details.
@@
    Demonstrating KeyMPs’ effectiveness in simulated environments through a comprehensive analysis of an occlusion-rich tasks, executable 3D motions that align with the intent expressed in the multimodal input, and showcasing its superior task generalization compared to deep learning-based and ablation methods.
@@ II-A Deep Learning-based DMPs Frameworks with Vision Input
Among these deep-learning-based DMP frameworks, only CIMEDNet [14], which generates motion by using a set of DMP parameters, and our prior work on DSDNet [21], which sequences DMPs by generating multiple sets of DMP parameters, are capable of planning complete task-oriented motion sequences from an image input. Despite this capability, the performance of both models remains severely limited by a lack of training data, which hinders their adaptability to new situations and tasks. To address these limitations, we used cross-domain knowledge through VLMs in this research to provide an understanding of visual input and essential insights to develop intricate robotic motions without requiring extra fine-tuning and that adapt to new situations.
@@ II-B LLM and VLM Integrated DMP Frameworks
Recent research has explored combining LLMs with DMPs in order to enhance robotic motion generation [19, 18]. It has shown that high-level language instructions can be translated into low-level motion primitives for complex task planning. However, as tasks become increasingly detailed or require contextual information from visual inputs, relying solely on language-driven approaches can make it challenging to generate precise positional data [30] or manage different groups of DMP parameters [19, 18].
@@ III-B Sequencing Multiple DMPs in Time
For more intricate behaviors, multiple DMPs can be sequenced in time [37, 38, 39, 40], with each handling a segment of the motion. Here, let $K$ be the total number of segments, with time boundaries $t_{0}\;<\;t_{1}\;<\;\cdots\;<\;t_{K},$, where $t_{0}$ and $t_{K}$ are the start and end times of the entire motion. The overall trajectory $\mathbf{Y}(t)$ is defined as a piecewise function:
@@ IV-C1 Keyword Labeled Primitive Selection
This approach requires collecting a small dataset of basis function weights for the specific primitives needed for the task. Since there are no universal rules for selecting primitives across all tasks, we therefore rely on domain knowledge to guide the selection process. For example, in the object cutting task, the required cutting primitives often vary depending on the object’s properties such as hardness or texture. Harder or crusty objects may require a sawing primitive, while softer objects can be cut using a downward primitive. Ultimately, the selection of appropriate primitives depends on the specific task at hand.
@@ IV-D DMP-based Motion Generation
Algorithm 1 KeyMPs Motion Generation
1:Parameters:
2:  $img_{env}$ Environment Visual Input
3:  $l$ Natural Language Input
4:  $h$ Object’s Height
5:Initialize:
6:  $D$ Reference Primitive Dictionary
[…]
11:return  $motion\ sequence$
@@ V SIMULATION EXPERIMENT
To evaluate the effectiveness of our KeyMPs framework and address key research questions, we selected two occlusion-rich tasks: object cutting task and cake icing task. Both tasks require one-shot complex motion generation because the knife or piping bag would occlude visual observations during execution. We created an environment in Isaac Gym [41] for both tasks, as shown in Fig. 5(a) for the object cutting task and Fig. 5(b) for the cake icing task. We conducted four experiments designed to test various aspects of our framework using these two tasks.
Figure 5: Simulation environments in Isaac Gym for (a) the object cutting task and (b) the cake icing task. In both scenarios, the scene is observed by a single top-down RGB camera. During execution, the robot’s tool obstructs the view of the object, which makes a one-shot motion generation approach essential.
@@ V-A Research Questions
  • (P4)
    Can the proposed method demonstrate generalizability to tasks other than cutting? (§V-G)
In the first experiment, we compared methods that directly generate DMP-based motion using VLMs to address (P1). For the second experiment, we checked the contribution of each VLM-based component separately. Finally, for the third experiment, we tested the generalization performance of our method by comparing it with existing deep-learning-based DMP frameworks. Finally, for the fourth experiment, we assessed the VLMs’ ability to generate complex keypoint pairs for cake icing tasks.
@@ V-B Experimental Setting
We performed three experiments focusing on the object cutting task to answer our research questions. Below, we outline the task details that serve as the foundation for our investigation.
V-B1 Object Cutting Task
Task Description
@@ V-B1 Object Cutting Task
Trained Task
1 Cabbage A single horizontal slice in the middle
@@ 9 Eggplant The object is 30 cm long, cut it vertically into 5 cm slices
Unseen Task
10 Cabbage Slice the object into 3 parts horizontally
@@ 15 Baguette The object is 45 cm long, cut it vertically into 5 cm slices
Evaluation Method
For each scenario in Table I, we first established ground-truth trajectories by applying kitchen-based domain knowledge to convert textual commands into precise geometric paths, such as by measuring an object’s dimensions to calculate a series of evenly spaced cuts. To validate that these patterns represent a common consensus, we subsequently conducted a user survey with 5 subjects (see Appendix E). The survey confirmed that our predefined lines aligned with the most widely agreed-upon cutting strategies for the given tasks. These validated 2D patterns were then systematically converted into 3D ground-truth coordinates by scaling the relevant primitive with the object’s height.
These 3D coordinates served two purposes. First, they were the benchmark for our quantitative evaluation. Second, for Experiment 3, they were used to generate the training data for the deep-learning baselines through imitation. To ensure the evaluation captured robustness, these ground-truth motions were also applied to objects with slightly randomized dimensions, accounting for subtle variations in object size.
Figure 6: Evaluation visualization comparing generated motion (light blue) to ground-truth motion (red). Each unpaired ground-truth point (crimson) is matched to the closest unpaired point in the generated motion (blue). This metric addresses variability in the number of points and focuses the comparison on the significant parts of the trajectory, excluding preparation or finishing motions.
For a quantitative evaluation (as shown in Fig. 6), a standard Mean Squared Error (MSE) calculation is not ideal due to potential variations in the number of points between the generated and ground-truth trajectories, as well as the presence of preparation or finishing motions that are not critical to the core cutting action. To address these issues, the ground-truth motion was first downsampled to a fixed number of points to ensure a fair comparison. Then, for each point in the ground truth trajectory, we find the index $j_{i}$ of the closest unpaired point in the generated trajectory and calculate the Euclidean distance $d_{i}$ between them:
$d_{i}=\|\mathbf{m}_{i}-\mathbf{n}_{j_{i}}\|$ (12)
where $i=1,2,\dots,N$, with $N$ being the number of points in the ground truth trajectory, $\mathbf{m}_{i}$ is the $i$-th point in the ground truth trajectory, $\mathbf{n}_{j_{i}}$ is the closest unpaired point in the generated trajectory to $\mathbf{m}_{i}$, and $j_{i}$ is the index of that point.
The overall evaluation metric $D$ is computed as:
$D=\frac{1}{N}\sum_{i=1}^{N}d_{i}$ (13)
This metric $D$ provides a quantitative measure of the discrepancy between the ground truth trajectory and the generated trajectory, accounting for spatial accuracy and ensuring that each ground truth point is uniquely matched to a generated point. A lower value of $D$ indicates a higher similarity between the generated motion and the ground truth, reflecting better performance of the method. This process enabled a consistent and accurate comparison between the generated and intended motions, focusing the comparison on the significant parts of the trajectory, excluding preparation or finishing motions.
It is important to note that this metric evaluates geometric proximity and is therefore insensitive to the temporal sequence of the trajectory. An otherwise accurate set of points executed in an illogical or scrambled order could still yield a low error score, as path coherence is not assessed.
Primitive Dictionary Preparation
We prepared two cutting primitives, selected from common motions in basic knife techniques [42], by having DMPs imitate predefined keypoints:
@@
V-B2 Cake Icing Task
Task Description
The objective of the cake icing task is to create icing designs on a cake according to the input prompt. The complexity of the designs depends on the intricacy of the desired outcome, impacting the generated keypoint pairs. Since the task is qualitative, we designed several cases shown in Table IV to assess the VLMs’ capabilities in generating intricate icing designs.
Evaluation Method
Given the qualitative nature of the cake icing task, evaluation was based on visual inspection and assessment of how well the generated designs aligned with the intended outcomes described in the input prompts. We examined whether the VLMs could produce complex keypoint pairs that resulted in creative and accurate icing designs, as specified in the task cases.
Primitive Dictionary Preparation
For the cake icing task, the framework uses a single icing line primitive [line] designed to emulate standard decorating techniques [43]. The primitive’s core motion involves a trajectory that first descends to the working surface. It then translates laterally to form a continuous line or decorative border. A visualization of this motion is available in Appendix D.
This single-primitive approach streamlines the framework. It places greater emphasis on the VLM’s ability to generate intricate keypoint designs. Its versatility is such that it can also produce icing dots (rosettes). This happens whenever the VLM specifies a keypoint pair with identical start and end coordinates.
@@ V-D Experiment 1: Comparison with Direct VLMs-to-DMPs Approach
Figure 12: Task-feature-extractor initialization prompt.
@@ V-D1 Objective
In this experiment, we evaluated the extent to which the structured component decomposition approach implemented in KeyMPs produces executable motions that align with the intent expressed in the multimodal input. Specifically, we compared KeyMPs against an unstructured, end-to-end VLM-driven approach that directly generates motion primitives. This comparison tested whether breaking the motion generation process into subtasks improves the capture of 3D geometric details and DMP parameters, resulting in motions that more faithfully reflect the intended outcome. Each task in Table I was executed ten times and was quantitatively evaluated using the criteria in §V-B1.
@@ V-E1 Objective
To address (P2), we evaluated the individual impacts of the two VLM-based components in our framework, i.e. keypoint pairs generation and keyword labeled primitive selection. This ablation study isolated each component’s effectiveness in producing executable motions aligned with the intended outcome. Each task in Table I was executed ten times and assessed using the criteria in §V-B1.
@@ V-F1 Objective
To address (P3), we determined whether KeyMPs achieves better generalization to unseen cutting tasks than deep-learning methods when training data is limited. The deep-learning baselines were trained on a restricted dataset of 90 demonstrations (10 per task for 9 tasks in Table I), simulating scenarios where collecting large-scale robotic data is impractical. In contrast, KeyMPs doesn’t require extensive task-specific training data, relying instead on structured component decomposition and vision-language priors. Each task in Table I was executed ten times, and the quantitative evaluation (§V-B1) focused on generalization performance, specifically success rates on novel objects and cutting patterns.